2 components combine to form 2 components. 0.1 mole of xyz weight 10 gm and 0.05 moles of xys weight 9 gm . calculate the atomic mass of xyz
1
Expert's answer
2019-06-26T04:23:52-0400

100 g/mol

Atomic mass of xyz is equal to A (xyz) = m (xyz) / n (xyz) = 10/0.1 = 100 g/mo